Jenny drove to her cabin on the lake and back. It took 0.6 hours longer to get there than it did to get back. The average speed of the trip was 46 km/h. The average speed of the trip back was 52 km/h. How many hours did the trip there take?
Answer by lwsshak3(6456)

Jenny drove to her cabin on the lake and back. It took 0.6 hours longer to get there than it did to get back. The average speed of the trip was 46 km/h. The average speed of the trip back was 52 km/h. How many hours did the trip there take?
**
let x=distance to cabin
travel time =distance/speed
x/46-x/52=0.6
LCD: 46*52
52x-46x=0.6*46*52
6x=1435.20
x=1435.20/6=239.2 km
Travel time to cabin=x/46=239.2/46=5.2 hrs
Travel time back=x/52=239.2/52=4.6 hrs
Difference=5.2-4.6=0.6 hrs